1. Solve for the area in square units of the right triangle with vertices F(0,7),U(2,-3),N(7,-2)




2. Find the area in square units of a rectangle whose vertices are L(0,6),O(2,-2),V(-2,-3),E(-4,5)






1
Expert's answer
2021-12-09T03:21:59-0500

1.

"\overrightarrow{UF}=\langle0-2, 7-(-3)\rangle=\langle-2,10)\rangle"

"\overrightarrow{UN}=\langle7-2, -2-(-3)\rangle=\langle5,1)\rangle"

"\overrightarrow{UF}\cdot\overrightarrow{UN}=-2(5)+10(1)=0=>\overrightarrow{UF}\perp\overrightarrow{UN}"

The right triangle "FNU"


"|\overrightarrow{UF}|=\sqrt{(-2)^2+(10)^2}=2\sqrt{26}"

"|\overrightarrow{UN}|=\sqrt{(5)^2+(1)^2}=\sqrt{26}"

"S=\dfrac{1}{2}|\overrightarrow{UF}||\overrightarrow{UN}|=\dfrac{1}{2}(2\sqrt{26})(\sqrt{26})=26({units}^2)"

2.


"\overrightarrow{VO}=\langle2-(-2), -2-(-3)\rangle=\langle4,1)\rangle"

"\overrightarrow{VE}=\langle-4-(-2), 5-(-3)\rangle=\langle-2,8)\rangle"

"\overrightarrow{EL}=\langle0-(-4), 6-5\rangle=\langle4,1)\rangle=\overrightarrow{VO}"

"\overrightarrow{OL}=\langle0-2, 6-(-2)\rangle=\langle-2,8)\rangle=\overrightarrow{VE}"




"\overrightarrow{VE}\cdot\overrightarrow{VO}=-2(4)+8(1)=0=>\overrightarrow{VE}\perp\overrightarrow{VO}"



The rectangle "VELO"


"|\overrightarrow{VE}|=\sqrt{(-2)^2+(8)^2}=2\sqrt{17}"

"|\overrightarrow{VO}|=\sqrt{(4)^2+(1)^2}=\sqrt{17}"

"S=|\overrightarrow{VE}||\overrightarrow{VO}|=2\sqrt{17}(\sqrt{17})=34({units}^2)"